ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

17. Control Theory and Optimization

1991 MS Classification: 65K10, 90C31, 90C33 Gomez Bofill, Walter, Mathematisch-Naturwissenschaftliche Fakultaet II, Institut fur Mathematik, Angewandte Mathematik, Humboldt Universitaet zu Berlin, D 10099 Berlin, Germany On the regularity of one-parametric variational Inequalities One parametric finite dimensional variational inequalities are considered. The regularity of one-parametric optimization problems in the sense of Jongen, Jonker and Twilt is formulated for the case of variational inequalities. Finally an algorithm for the resolution of variational inequalities is analysed with this approach.